Twenty four students are in Miss favor's class. One sixth 1/6 of the students were born in the first quarter of the year. One half(1/2) of the remaining students were born in the months May through august. There are no september birthday but one fifth(1/5) of the remaining birthdays are in october. If one student was born in november, how many december birthdays are there?
Answer by ikleyn(52782)

.
Twenty four students are in Miss favor's class.
One sixth 1/6 of the students were born in the first quarter of the year.
One half(1/2) of the remaining students were born in the months May through august.
There are no September birthday
but one fifth(1/5) of the remaining birthdays are in October.
If one student was born in November, how many December birthdays are there?
~~~~~~~~~~~~~~~~~~~~~

See the solution in the table below

    Months   n of birthdays   cumulative sum    n of remaining  
                               of birthdays        birthdays      
  --------------------------------------------------------------

    1-3       24/6 = 4            4              24-4 = 20
     
     4          0                 4                20

    5-8       (1/2)*20        4+10 = 14          24-14 = 10
               = 10

     9          0                14                10

    10        (1/5)*10        14+2 = 16          24-16 = 8
                = 2         

    11          1 (one)       16+1 = 17          24-17 = 7


ANSWER.  7 December birthdays.

Solved.